40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 19 14:06:36 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_14-06-11.png
    Reply  Fri Apr 19 20:21:52 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_20-27-04.png
       Reply  Wed Apr 24 16:06:44 2019, gautam, Update, PSL, Innolight NPRO shutoff PMCrefl.png
          Reply  Thu Apr 25 01:03:29 2019, gautam, Update, PSL, MC trans is ~1000 cts (~7%) lower than usual 
          Reply  Thu Apr 25 17:31:56 2019, gautam, Update, PSL, Innolight NPRO shutoff EdwinShutoff20190425.pngEdwinShutdown_zoomOut.png
             Reply  Mon Apr 29 16:25:22 2019, gautam, Update, PSL, PSL turned on again 
                Reply  Fri May 3 12:51:58 2019, gautam, Update, PSL, PSL turned on again 
                   Reply  Mon May 6 10:51:43 2019, gautam, Update, PSL, PSL turned off again NPROshutoff.png
                   Reply  Wed May 8 19:04:20 2019, rana, Update, PSL, PSL turned on again 
                      Reply  Thu May 9 19:50:04 2019, gautam, Update, PSL, PSL turned off again Screenshot_from_2019-05-09_19-49-29.png
                         Reply  Mon May 13 10:45:38 2019, gautam, Update, PSL, PSL turned ON again IMG_7431.JPG
                            Reply  Wed May 15 10:56:47 2019, gautam, Update, PSL, PSL turned ON again 
Message ID: 14605     Entry time: Mon May 13 10:45:38 2019     In reply to: 14599     Reply to this: 14609
Author: gautam 
Type: Update 
Category: PSL 
Subject: PSL turned ON again 

I used some double-sided tape to attach a San Ace 60 9S0612H4011 to the Innolight controller (Attachment #1). This particular fan is rated to run with up to 13.8V, but I'm using a +15V Sorensen output - at best, this shortens the lifespan of the fan, but I don't have a better solution for now. Then I turned the laser on again (~1040 local time), using the same settings Rana configured earlier in this thread. PMC was locked, and the IMC also could be locked but I closed the shutter for now while the laser frequency/intensity stabilizes after startup. The purpose is to facilitate completion of the pre-vent alignment checklist in prep for the planned vent tomorrow. PMC Trans reports 0.63 after alignment was optimized, which is ~15% lower than in Oct 2016.

Attachment 1: IMG_7431.JPG  1.004 MB  Uploaded Mon May 13 12:53:31 2019  | Hide | Hide all
IMG_7431.JPG
ELOG V3.1.3-